I have been working on a problem using the elimination method. I think I have it right but would someone tell me if it is and if not what I have done wrong. The problem is 5r - 2s = 8
2r + 5s = 38
I do not think there is a solution.

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

multiplying 1st eqn by 5 ___ 25r - 10s = 40

multiplying 2nd eqn by 2 ___ 4r + 10s = 76

adding new eqns (this eliminates the s terms) ___ 29r = 116 ___ r = 4

substituting ___ 2(4) + 5s = 38 ___ 5s = 30 ___ s = 6

Answer by stanbon(57387)

The problem is
5r - 2s = 8
2r + 5s = 38
-----------------------------
Multiply thru the 1st by 2
Multiply thru the 2nd by 5
-----------------------------
10r - 4s = 16
10r + 25s = 190
----------------------
Subtract the 1st Eq. from the 2nd Eq and solve for "s":
29s = 174
s = 6
--------------
substitute into 5r-2s=8 to solve for "r":
5r -2*6 = 8
5r = 20
r = 4
--------------
Solution: (4,6)
=======================
Cheers,
Stan H.